OpenStudy (anonymous):

A drought can cause the level of the local water supply to drop by a few inches each week.Suppose the level of the water supply drops 2.8 inches each week A. how far will it have dropped in 4 weeks. B. How far will it have dropped in 6 half weeks C. Create another real world situation which is not a water supply drought,using the numbers -2.8 and 4. its multiplying rational numbers

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@mathaholic

OpenStudy (anonymous):

@TuringTest

OpenStudy (mathstudent55):

If the water level drops 2.8 inches in 1 week, then you can construct the table below: Week Water drop 1 2.8 = 1 * 2.8 = 2.8 2 2.8 + 2.8 = 2 * 2.8 = 5.6 3 5.6 + 2.8 = 2.8 + 2.8 + 2.8 = 3 * 2.8 = 8.4 4 8.4 + 2.8 = 2.8 + 2.8 + 2.8 + 2.8 = 4 * 2.8 = 11.2 You see that part A is a multiplication problem. The water level drops 2.8 inches in 1 week, so to find how much it drops in any number of weeks, multiply the number of weeks by 2.8.

OpenStudy (mathstudent55):

For part B, all you need to do is multiply the number of weeks by the drop, 2.8. First, how many weeks is 6 half weeks? Figure that out then multiply it by 2.8.

OpenStudy (anonymous):

What about C

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Euler271